OpenStudy (anonymous):

What is the x-intercept of the line with the equation 1/3x + y = -15

OpenStudy (dayakar):

to find x-intercept write y=0 in the equation

OpenStudy (dayakar):

can u do it

OpenStudy (dayakar):

1/3x +0 = -15 1/3x= -15

OpenStudy (dayakar):

x = -15 *3/1 x= -45 x-intercept = -45

OpenStudy (dayakar):

get it